If you are looking at a graph of a quadratic equation y = ax2 + bx + c, how do you determine where the solutions to 0 = ax2 + bx + c are? Compare this with solving the equation using the quadratic formula.
Answer by stanbon(75887)

If you are looking at a graph of a quadratic equation y = ax2 + bx + c, how do you determine where the solutions to 0 = ax2 + bx + c are? Compare this with solving the equation using the quadratic formula.
----------
Formula: x = [-b +- sqrt(b^2-4ac)]/(2a)
----
If b^2-4ac = 0, solutions are -b/2a and are two equal Real Numbers
If b^2-4ac < 0, solutions are two complex numbers.
If b^2-4ac > 0, solutions are 2 unequal Real Numbers.
=================
Cheers,
Stan H.
=================